Shift Summary 30 June Day shift.

Nickolas Solomey (solomey@hep.uchicago.edu)
Wed, 30 Jun 1999 15:27:43 -0500

Shift Summary for day shift, 29 Jun 99

On Shift: Nick and Jianbo

Hours of beam delivered at all intensities: 6.25
Hours of beam delivered at nominal intensity: 6.25 (5.1e12)
Hours of E832 data (special runs not included): 6.25
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: 65 min Tevatron problems, 55 min. shutting
down computers and changing a tape drive and rebooting the computer,
plus at least another 15 min of single spills lost due to SY energy
conservation.

========================================================================
Runs ended this shift:
Run 13772
Run 13773
Run 13774

Tapes brought to FCC this shift: None

========================================================================
Accesses during shift: Yes, to change an ET board.

========================================================================
Overview of shift: A good fraction of E832 data taking, but
some excitment when the computer division came over to change our
tape drive.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13772 160 07:00 Take over rum from previous shift.
169 7:14 Found 40 hotbits at chan 662
180 7:30 The count of hotbits at chan 662 goes up to 116
185 Trigger in error state, did a kfrend
186 No beam that spill
221 8:23 The count of hotbits at chan 662 goes up to 1145
230 8:36 Beam vanishes
8:45 MCR> Tev quad corrector repair underway,
estimate 15 min.
243 8:50 Beam returns
257 9:11 Turned off Expresso machine Jason L. left on.
268 9:27 HCC inefficiency, Chan 662 and 248. The count of
hotbits at chan 662 goes up to 2198.
10:12 MCR> Fridge permit lost, estimate 10 min.
for beam to return.
318 10:32 Beam returns
13772 329 10:48 last spill, run ended for computer repair.
11:02 JH and JBW make access to swap ET board.
11:12 Return from acess
11:48-52 Dave N. and Hogan made access to get some
DC equipment and Hogan checked DC2 banna cables
for Jason L.
12:00 Called in to MCR we are out of access and ready
to run.
13773 Run started, but ended without data to update
Fera pedistals.
13774 1 12:30 Beam_832 run started.
12:46 Paged Sydney because of onsplit error, he fixed
by phone.
31 13:10 Ktevana error, ordinal number mismatch, run
kfrend.
53 13:38 Laser stability to wide, but went away the next
spill.
136 15:30 End of shift, run continues into the next shift.